OpenStudy (anonymous):

Please help! Write the equation of a parabola with a vertex at (5, 2) and a directrix y = 1.

OpenStudy (anonymous):

x-5=1/4(y-2)^2 y-2=1/4(x-5)^2 x-2=1/4(y-5)^2 y+2=1/4(x-5)^2

OpenStudy (welshfella):

use the general form (x - h)^2 = 4p(y - k) where (h,k) is the vertex and y = p is the directrix plug in the values of h,k and p